OpenStudy (anonymous):

how do i solve the polynomial equation x^2+9x+5=x-5

OpenStudy (anonymous):

First you must set the equation equal to 0, so the x must be subtracted on each side: x^2+9x+5=x-5 turns to x^2+8x+5=-5 Now to get rid of the pesky -5: x^2+8x+5=-5 turns to x^2+8x+10=0 Remember when there's a negative you want to rid of, you add to both sides. When there's a positive you want to rid of, you subtract to both sides. Now x^2+8x+10 cannot be factored out well because no numbers will add to +8 and multiply to +10 so you use the formula: [-b±√(b^2-4ac)]/2a Now, using x^2+8x+10 as a form of y=ax^2+bx+c, you can determine: 1=a 8=b 10=c Now plug in the numbers and solve for the equation.

OpenStudy (anonymous):

is the answer -8+- square root of 24/2?

OpenStudy (anonymous):

(-8±√24)/2 From my calculations, yes. I am not absolutely sure if this is the correct answer but I've worked it out thrice and came with the same answer.

OpenStudy (anonymous):

thankyou.